Solve and give solution set for the following equation in R

(a) 2X+3<3X+4>X+5

(b) 2x^2+4=0

(c) 2sin ( 180-x) =1

(d) (X^2-X-6) ÷ (X^2-X-2)<0


1
Expert's answer
2021-09-03T16:09:24-0400

(a)


"2x+3<3x+4>x+5"

Then


"\begin{matrix}
 2x+3<3x+4 \\
 3x+4>x+5
\end{matrix}"

"\begin{matrix}
 x>-1 \\
 2x>1
\end{matrix}"

Answer: "x>\dfrac{1}{2}"


"x\in(\dfrac{1}{2}, \infin)"


(b)

"2x^2+4=0"

"x^2\geq0, x\in\R"

"2x^2+4\geq4, x\in\R"

The given equation has no solution for "x\in\R."



"2x^2+4=0"

"x^2=-2"

"x_1=-i\sqrt{2}, x_2=i\sqrt{2}"

(c)


"2\sin(180-x) =1"

"2\sin x=1"

"\sin x=\dfrac{1}{2}"

"x=(-1)^n\cdot30\degree+180\degree n, n\in\Z"

(d)


"\dfrac{x^2-x-6}{x^2-x-2}<0"




"(x^2-x-6)(x^2-x-2)<0"


"(x+2)(x-3)(x+1)(x-2)<0"


"x\in(-2, -1)\cup(2, 3)"